# Targeted Nakamoto: A Bitcoin Protocol to Balance Network Security and Energy Consumption

### v3

@@ 3.2 An overview of the mechanism design of the protocols
Targeted can be decomposed into two protocols that interact while operating separately. The Nakamoto part is the mechanism associated with the current protocol. The Targeted part adds a mechanism to achieve a target hahsrate interval. Figure 3 displays the information recorded on the blockchain that is used by each protocol [^17] The network also records the addresses of UTXO’a at each block. Neither Nakamoto or Targeted use that information..
@@ 4.1 The puzzle difficulty signal of hashrate
A single mining computer can guess a puzzle solution sequentially in time. [^19] A computer with $n$ parallel processors is treated here as $n$ separate computers. I model mining computers as making successive guesses synchronously at times $t\in\{t_{1},t_{2},...\}$, each separated by an interval of temporal length $t_{i+1}-t_{i}$ which I normalize to 1 unit of time [^20] The synchronicity enables the expression of the time-path of puzzle solutions compactly as a Bernoulli Process, which simplifies the analysis WLOG.. The Nakamoto blockchain growth target is one block every $T^{*}$ units of time. During this time interval each computer has guessed (or hashed) $T^{*}$ proposed solutions. I define a unit of hashrate as a single mining computer making $T^{*}$ puzzle guesses. If there are $N$ mining computers making guesses, the number of total guesses made in the time interval of the blockchain growth target is $NT^{*}$ and the hashrate is $N$212121 This accords with the conventional definition of hashrate used e.g. by Coinmetrics [4], as puzzle guesses per second. The framework for guessing puzzle solutions over time is a Bernoulli Process. Let $T$ denote the time the first puzzle solution is broadcast to the network. For a hashrate of $N$ the expected time to reach the solution is expressed by Equation 2.
[…]
Equation 2 shows that, in order to achieve the growth rate target of $T^{*}$ in expectation, puzzle difficulty $D$ must be adjusted to offset hashrate $N$. Since $T$ for the prior epoch is recorded on the blockchain and $D$ is set by the code, the formula can be inverted to estimate hashrate as a function of the growth rate and puzzle difficulty, $\hat{N}=D/T$232323 $\hat{N}=N+\epsilon$, where $\epsilon$ is a mean zero iid random variable. $\hat{N}$ is unbiased since the observed $T$ is an unbiased estimate of the mean time interval implied by the Bernoulli Process. It is also a sufficient statistic for hashrate since $D$ and $T$ contain all of the information on $N$ that is encoded on the blockchain.
@@ 4.2 Mining equilibrium
The miner who is first to broadcast a puzzle solution receives a block reward of cryptocurrency. Miners incur cost in terms of the local official currency, which we denote USD. [^24] Profit can be expressed in terms of cryptocurrency. In that case the unit hashrate cost cost $c$ is multiplied by $1/e$, which is the USD to BTC exchange rate. The equilibrium hashrate is the same whatever currency is chosen to be the unit of account. Mining equilibrium is a state where miners choose hashrate to maximize expected profit. When miners adjust hashrate at short block intervals profit maximization implies that the profit from a marginal unit of hashrate equals the competitive rate of profit. The independence of puzzle guesses means each unit of hashrate has a $1/N$ probability of receiving the block reward. Equation 3 is the equilibrium profit from mining one unit of hashrate (i.e. operating one mining computer for $T^{*}$ units of time). [^25] Equation 3 is a dynamic equilibrium in the sense that it includes a growth rate that deviates from target. A rigorous expression would show $\Delta_{T}$ as a function of puzzle difficulty and hashrate. We do not derive the equilibrium from fundamentals and therefore we do not prove the stability of the solution.
@@
$\Delta_{T}=T^{*}-T$ is the difference between the target growth rate, $T^{*}$, and the expected actual growth rate $T$ [^26] Blockchain growth rate is expressed as an expectation since it is stochastic. $(1+\Delta_{T})$ is the number of blocks per $T^{*}$ units of time; $c$ is the marginal cost of a unit of hashrate (i.e. one computer for $T^{*}$ units of time); [^27] Marginal cost is the cost of electricity to run the ASIC’s. Fixed costs, such as the rent on the building where the computers are stored, or the sunk cost of purchasing the computers, are not included. $e$ is the USD/BTC exchange rate and $p_{B_{n}}^{M}$ is the expected miner’s block reward on the $n^{th}$ block. ($p_{B_{n}}$ is the total block reward). I make no assumption about the structure of the mining market. $\phi\geq 0$ a competition parameter. Under free entry in the mining market $\phi=0$. Under less competitive conditions $\phi>0$. Equation 4 is the mining market equilibrium.
@@
Equation 5 displays the monotone relationship between the miner’s block reward, $p^{M}_{B}$, and the sufficient statistic for hashrate, $D/T$. The dynamic adjustment of hashrate to changes in the block reward is evaluated in Section 6.
@@ 5.3 Policy to control hashrate
For each subsequent epoch until $\text{floor}_{q}\geq\overline{P_{B}}(q)$, the floor is reduced by the formula $\text{floor}_{q+1}\leftarrow\gamma\text{floor}_{q}$ Thereafter, the floor is removed.
 [^31] The formulas $\tau$ and $(1+(1-\tau))$ ensure symmetry in the adjustment of the ceiling and the floor, i.e. that the ceiling and floor adjust by the same percentage each epoch.
Finally, note that the Hashrate Control Algorithm is a hard-fork, since the imposition of the floor to the miner’s block reward can require the minting of coinbase that exceeds the amount scheduled in Nakamoto.
6 Dynamic Adjustment of Hashrate
@@
Consider an epoch $q$ where the total block reward is constant at $p^{M}_{B}$ and the puzzle difficulty adjustment at the end of epoch $q-1$ achieves the target growth rate when hashrate is $N$, and $N$ is the equilibrium hashrate when the miner’s block reward is $p^{M}_{B}$. Suppose a ceiling or floor is imposed at the end of epoch $q-1$ resulting in a marginal change in the miner’s block reward equal to $\Delta p$ (i.e. holding fixed the exchange rate, $e$, and hashrate cost $c$). Let $\Delta_{T}$ denote the distance between the actual blockchain growth rate and the target blockchain growth rate induced by a change in hashrate in epoch $q$, denoted $\Delta N$.
@@
[…]
The change in hashrate can be decomposed into a lottery effect and a growth rate effect. The lottery effect reflects that a change in hashrate changes the odds of any unit of hashrate mining the next block. The growth rate effect reflects that a change in hashrate alters the blockchain growth rate, which changes the number of puzzle lotteries in which a unit of hashrate can compete in $T^{*}$ units of time. Noting that $\frac{D}{N}|modulo\;T^{*}=1$, the growth rate effect is $2-\frac{D}{N+\Delta N}$. Equation 7 is the marginal effect of hashrate on profit.
@@
[…]
@@ 7.1 The Targeted Monetary Policy
Additions to the total block reward: When $q$ is added to the total block reward at block $B_{n}$ there are two possible adjustments to accounts. If $P-q>0$, $x$’s assigned value in the UTXO pool is reduced by the product of its percentage times $q$, rounded down to the nearest Satoshi. When $P-q\leq 0$ two accounting changes occur; (i) P and $x$’s assigned value $e_{x}$ are each reduced to zero and (ii) the value $q-P$ is added to an account that tracks value that is subtracted from the spending potential of addresses (the ”Remainder Accounts”) which has initial value $R$. Each address $x$ is assigned a value in the remainder account, denoted $r_{x}$, which is increased by its percentage times $P-q$, rounded down to the nearest Satoshi. [^37] The UTXO value subtracted from P can either be burned or sent to the miner. The latter would reduce the minting of UTXO required to increase the miner’s block reward.
@@
Figure 5 depicts the operation of the Targeted monetary policy at block $B_{n}$. The timing sequence is denoted by the circled numbers. The first step is the execution of the transactions. The adjustment to a transaction sender’s UTXO value is determined by the difference between its account value in the UTXO pool and its account value in the Remainder Accounts. The second step is the payment of the miner block reward. If $P^{M}_{B_{n}}<p^{M}_{B_{n}}$, UTXO value of $q$ is minted to cover the payment. The third step is to send of UTXO value $q$ to the UTXO Pool if $P^{M}_{B_{n}}<p^{M}_{B_{n}}$ or else burn UTXO value up to $q$ in the UTXO pool and increase the value of the Remainder Accounts by the residual. Adjustments to the address account values in the UTXO Pool and the Remainder Accounts are made.393939 Steps 2 and 3 can be simultaneous. When $q<0$ the minting of UTXO value is reduced by the amount of value in the UTXO pool, which is sent to the miner from the UTXO pool. This is implicit in the code in Appendix B.
@@ 8 Conclusion
Targeted is motivated by two observations about PoW blockchains. One observation is that there are costs which increase with hashrate - electricity and carbon emissions - and costs which decrease with hashrate - network security. This observation suggests there is an interval of hashrate in which total costs are minimized. The other observation is that hashrate is a function of the value of the block reward, which is unbounded. The Nakamoto protocol does not provide an incentive for miners to avoid hashrate dropping to a very low level or reaching a very high level. Targeted incentivizes miners to home in on a target hashrate interval. This is achieved by imposing a floor to the miner’s block reward when hashrate is below target and imposing a ceiling on the miner’s block reward when hashrate is above target. The ceiling is progressively lowered and the floor is progressively raised for so long as hashrate is outside the target interval. Targeted achieves aggregate and distributional monetary neutrality. Aggregate monetary neutrality means that the UTXO spending potential of the network is the same under Targeted and Nakamoto at each block. Distributional monetary neutrality means that the UTXO spending potential of addresses are adjusted to offset increases and decreases in the block reward in proportion to their UTXO value.
There are two preconditions for implementing Targeted. One is that a target electricity consumption interval needs to be identified. The second is that the efficiency of mining computers needs to be estimated and recalibrated over time in order to adjust the target hashrate interval to ensure it maps into the target electricity consumption interval. For example, when ASIC’s become more efficient, the target hashrate interval will need to be adjusted upward, since each hash consumes less electricity.
Finally, I argued that there is an urgency to augment the Nakamoto protocol to control hashrate. The current perceived high level of electricity consumed by Bitcoin miners has contributed global warming and a political backlash against Bitcoin. This is a near term challenge for Bitcoin. In the long run Bitcoin faces an existential threat from the opposite direction. The scheduled halving of coinbase approximately every four years will eventually make transaction fees the primary source of miner revenue. Historically, transaction fees have comprised a small percentage of overall block reward. Therefore, in the absence of an orders of magnitude increase in transaction fees, hashrate and electricity consumption may decline to a low level that threatens network security.

### v4

@@ 3.2 An overview of the mechanism design of the protocols
Targeted can be decomposed into two protocols that interact while operating separately. The Nakamoto part is the mechanism associated with the current protocol. The Targeted part adds a mechanism to achieve a target hashrate interval. Figure 3 displays the information recorded on the blockchain that is used by each protocol [^14] The network also records the addresses of UTXO’a at each block. Neither Nakamoto or Targeted use that information..
@@ 4.1 The puzzle difficulty signal of hashrate
A single mining computer can guess a puzzle solution sequentially in time. [^16] A computer with $n$ parallel processors is treated here as $n$ separate computers. I model mining computers as making successive guesses synchronously at times $t\in\{t_{1},t_{2},$...\}, each separated by an interval of temporal length $t_{i+1}-t_{i}$ which I normalize to 1 unit of time [^17] The synchronicity enables the expression of the time-path of puzzle solutions compactly as a Bernoulli Process, which simplifies the analysis WLOG.. The Nakamoto blockchain growth target is one block every $T^{*}$ units of time. During this time interval each computer has guessed (or hashed) $T^{*}$ proposed solutions. I define a unit of hashrate as a single mining computer making $T^{*}$ puzzle guesses. If there are $N$ mining computers making guesses, the number of total guesses made in the time interval of the blockchain growth target is $NT^{*}$ and the hashrate is NN.181818 This accords with the conventional definition of hashrate used e.g. by Coinmetrics [5], as puzzle guesses per second. The framework for guessing puzzle solutions over time is a Bernoulli Process. Let $T$ denote the time the first puzzle solution is broadcast to the network. For a hashrate of $N$ the expected time to reach the solution is expressed by Equation 3.
$\underset{\text{expected time between blocks}}{\underbrace{\mathbb{E}[T]}}=D/N$ (3)
Equation 3 shows that, in order to achieve the growth rate target of $T^{*}$ in expectation, puzzle difficulty $D$ must be adjusted to offset hashrate $N$. Since $T$ for the prior epoch is recorded on the blockchain and $D$ is set by the code, the formula can be inverted to estimate hashrate as a function of the growth rate and puzzle difficulty, $\hat{N}=D/T$.202020 $\hat{N}=N+\epsilon$, where $\epsilon$ is a mean zero iid random variable. $\hat{N}$ is unbiased since the observed $T$ is an unbiased estimate of the mean time interval implied by the Bernoulli Process. It is also a sufficient statistic for hashrate since $D$ and $T$ contain all of the information on $N$ that is encoded on the blockchain.
@@ 4.2 Mining equilibrium
The miner who is first to broadcast a puzzle solution receives a block reward of cryptocurrency. Miners incur cost in terms of the local official currency, which we denote USD. [^21] Profit can be expressed in terms of cryptocurrency. In that case the unit hashrate cost cost $c$ is multiplied by $1/e$, which is the USD to BTC exchange rate. The equilibrium hashrate is the same whatever currency is chosen to be the unit of account. Mining equilibrium is a state where miners choose hashrate to maximize expected profit. When miners adjust hashrate at short block intervals profit maximization implies that the profit from a marginal unit of hashrate equals the competitive rate of profit. The independence of puzzle guesses means each unit of hashrate has a $1/N$ probability of receiving the block reward. Equation 4 is the equilibrium profit from mining one unit of hashrate (i.e. operating one mining computer for $T^{*}$ units of time). [^22] Equation 4 is a dynamic equilibrium in the sense that it includes a growth rate that deviates from target. A rigorous expression would show $\Delta_{T}$ as a function of puzzle difficulty and hashrate.
@@
$\Delta_{T}=T^{*}-T$ is the difference between the target growth rate, $T^{*}$, and the expected actual growth rate TT. [^23] Blockchain growth rate is expressed as an expectation since it is stochastic. $(1+\Delta_{T})$ is the number of blocks per $T^{*}$ units of time; $c$ is the marginal cost of a unit of hashrate (i.e. one computer for $T^{*}$ units of time); [^24] Marginal cost is the cost of electricity to run the ASIC’s. Fixed costs, such as the rent on the building where the computers are stored, or the sunk cost of purchasing the computers, are not included. $e$ is the USD/BTC exchange rate and $p_{B_{n}}^{M}$ is the expected miner’s block reward on the $n^{th}$ block. ($p_{B_{n}}$ is the total block reward). I make no assumption about the structure of the mining market. $\phi\geq 0$ a competition parameter. Under free entry in the mining market $\phi=0$. Under less competitive conditions $\phi>0$. Equation 5 is the mining market equilibrium.
@@
Equation 6 displays the monotone relationship between the miner’s block reward, $p^{M}_{B}$, and the sufficient statistic for hashrate, $D/T$.
@@ 5.3 Policy to control hashrate
For each subsequent epoch until $\text{floor}_{q}\geq\overline{P_{B}}(q)$, the floor is reduced by the formula $\text{floor}_{q+1}\leftarrow\gamma\text{floor}_{q}$ Thereafter, the floor is removed. [^28] The formulas $\tau$ and $(1+(1-\tau))$ ensure symmetry in the adjustment of the ceiling and the floor, i.e. that the ceiling and floor adjust by the same percentage each epoch.
6 Strategic Effects of Block Adjustment Policy
The adjustment of the miner’s block reward introduces a new variable into equilibrium profit formula of Equation 4. Treating the target hashrate as a point, $N^{T}=N_{UB}=N_{LB}$, the block reward can be expressed as $p^{M}_{B_{n}}(N^{T}-N)$. This shows that the block reward adjusts to deviations from the target hashrate. A question arises as to whether the block adjustment policy can induce strategic behavior that upsets the monotone response to block reward adjustments that was assumed in the previous section. Here we evaluate the stability of equilibrium at the targeted hashrate from two angles. First, we derive bounds on the block reward adjustment that ensure deviation from the equilibrium is unprofitable. Second, we show the adjustment mechanism causes hashrate to home in on the target under an assumption of constant miner profit. There may be other evaluations that could be carried out by placing different restrictions, but I believe the two chosen here show the reasonableness of the assumption that hashrate responds monotonically to changes in the block reward. [^29] A maintained assumption of this paper is that block timestamps are correct. I do not explore whether Targeted creates a new vulnerability to attack by miner manipulation of timestamps.
6.1 Stability of Miner Equilibrium
Equation 7 is the profit of a miner with hashrate $S$ at the target hashrate.
$\pi(S)=\frac{S}{N^{T}}eP(N^{T})^{M}_{B}(1-\Delta_{T})-cS$ (7)
The profit from deviating and adding or subtracting hashrate by $\Delta_{S}$ is
$\pi(S+\Delta_{S})=\frac{S+\Delta_{S}}{N^{T}+\Delta_{S}}eP(N^{T}+\Delta_{S})^{M}_{B}(1-\Delta_{T})-c(S+\Delta_{S})$
A deviation is unprofitable if $\pi(S+\Delta_{S})<\pi(S),\;\forall\Delta_{S}$. Equation 8 is a sufficient condition.
$P(N^{T}+\Delta_{S})<\frac{N^{T}+\Delta_{S}}{S+\Delta_{S}}\times\frac{S}{N^{T}}P(N^{T})^{M}_{B}$ (8)
Since $S\leq N^{T}$ by definition, it is immediate that, in order to maintain the inequality, an increase in hashrate, $\Delta_{S}>0$ must cause the block reward to decrease and a decrease in hashrate does the opposite. This is consistent with the block adjustment policy. We can derive bounds on the magnitude of adjustment by evaluating the required adjustment at the extremes of $\Delta_{S}=1$ and $\Delta_{S}=0$. This yields the range of adjustment multiple $(\frac{1+\Delta_{S}/N^{T}}{1+\Delta_{S}},1)$. Since the lower bond approaches 1 as $\Delta_{S}$ increases from 1 to $N^{T}$, the lower bound can be restricted to its value at $\Delta_{S}=1$, which yields $\frac{1+1/N^{T}}{2}$. [^30] It is straightforward to add this constraint to the block adjustment formula replacing the point $N^{T}$ with the interval interval $N_{LB},N_{UB}$.
6.2 Dynamic Adjustment of Hashrate
@@
Consider an epoch $q$ where the total block reward is constant at $p^{M}_{B}$ and the puzzle difficulty adjustment at the end of epoch q1q-1 achieves the target growth rate when hashrate is $N$, and $N$ is the equilibrium hashrate when the miner’s block reward is $p^{M}_{B}$. Suppose a ceiling or floor is imposed at the end of epoch q1q-1 resulting in a marginal change in the miner’s block reward equal to $\Delta p$ (i.e. holding fixed the exchange rate, $e$, and hashrate cost $c$). Let $\Delta_{T}$ denote the distance between the actual blockchain growth rate and the target blockchain growth rate induced by a change in hashrate in epoch $q$, denoted $\Delta N$.
@@
$\pi_{\Delta p,\Delta N}=\\ \underset{\text{lottery effect}}{\underbrace{\frac{1}{N+\Delta N}}}\Big{(}\underset{\text{growth rate effect}}{\underbrace{[1+\underset{\Delta_{T}}{\underbrace{D\{\frac{1}{N}-\frac{1}{N+\Delta N}\}}}}}](e(P_{B}^{M}+\Delta p)\Big{)}\\ -c=\phi$ (9)
The change in miner profit can be decomposed into a lottery effect and a growth rate effect. The lottery effect reflects that a change in hashrate changes the odds of any unit of hashrate mining the next block. The growth rate effect reflects that a change in hashrate alters the blockchain growth rate, which changes the number of puzzle lotteries in which a unit of hashrate can compete in $T^{*}$ units of time. Noting that $\frac{D}{N}|modulo\;T^{*}=1$, the growth rate effect is $2-\frac{D}{N+\Delta N}$. Equation 10 is the marginal effect of hashrate on profit.
@@
Using the implicit function theorem to combine equations 10 and 11 shows that an increase in the block reward induces an increase in hashrate (and vice versa) when hashrate adjusts to maintain a constant profit rate per unit of time $\Delta N/\Delta p=-\frac{\Delta\pi}{\Delta p}\big{/}\frac{\Delta\pi}{\Delta N}=+/+>0$. [^33] The analysis of dynamic equilibrium can be extended to the case where the difficulty adjustment at the end of epoch $q$ does not achieve the target growth rate. In that case the ratio $D/N\neq T^{*}$ and $\frac{\Delta\pi}{\Delta N}=-e(P_{B}^{M}+\Delta p)\frac{1}{(N+\Delta N)^{2}}\big{(}1+D/N\big{)}<0$ which yields the same result in terms of the direction of the response of hashrate to the value of the miner’s block reward. The key point is that, when hashrate is responsive to changes in the block reward, the marginal effect of the Hashrate Control Algorithm is to induce an adjustment of hashrate in the intended direction when a ceiling or floor pushes the miner’s block reward below or above the total block reward. [^34] Assuming that $e$ and $c$ are bounded, the result can easily be extended to show that the Hashrate Control Algorithm will eventually push hashrate toward the target interval, but it will not necessarily do so monotonically. The foregoing proves the following proposition. [^35] Bissias et.al. (2022) [2] document that miner spending allocation between Bitcoin and Bitcoin Cash adjust at high frequency to maintain a zero arbitrage condition. This validates that hashrate tends to adjust quickly. A richer formulation of my model would include the shadow profit of mining an alternative blockchain.
@@ 7.1 The Targeted Monetary Policy
Additions to the total block reward: When $q$ is added to the total block reward at block $B_{n}$ there are two possible adjustments to accounts. If Pq>0P-q>0, $x$’s assigned value in the UTXO pool is reduced by the product of its percentage times $q$, rounded down to the nearest Satoshi. When $P-q\leq 0$ two accounting changes occur; (i) P and $x$’s assigned value $e_{x}$ are each reduced to zero and (ii) the value qPq-P is added to an account that tracks value that is subtracted from the spending potential of addresses (the ”Remainder Accounts”) which has initial value $R$. Each address $x$ is assigned a value in the remainder account, denoted $r_{x}$, which is increased by its percentage times PqP-q, rounded down to the nearest Satoshi. [^36] The UTXO value subtracted from P can either be burned or sent to the miner. The latter would reduce the minting of UTXO required to increase the miner’s block reward.
@@
Figure 5 depicts the operation of the Targeted monetary policy at block $B_{n}$. The timing sequence is denoted by the circled numbers. The first step is the execution of the transactions. The adjustment to a transaction sender’s UTXO value is determined by the difference between its account value in the UTXO pool and its account value in the Remainder Accounts. The second step is the payment of the miner block reward. If $P^{M}_{B_{n}}<p^{M}_{B_{n}}$, UTXO value of $q$ is minted to cover the payment. The third step is to send of UTXO value $q$ to the UTXO Pool if $P^{M}_{B_{n}}<p^{M}_{B_{n}}$ or else burn UTXO value up to $q$ in the UTXO pool and increase the value of the Remainder Accounts by the residual. Adjustments to the address account values in the UTXO Pool and the Remainder Accounts are made.383838 Steps 2 and 3 can be simultaneous. When $q<0$ the minting of UTXO value is reduced by the amount of value in the UTXO pool, which is sent to the miner from the UTXO pool. This is implicit in the code in Appendix B.
@@ 8 Conclusion
Targeted is motivated by two observations about PoW blockchains. One observation is the two external hashrate costs of network security and $CO^{2}$ emissions, the former decreasing and the latter increasing in hashrate. This observation suggests there is an interval of hashrate in which total costs are minimized. The second observation is that hashrate is a function of the value of the block reward. Nakamoto does not provide an incentive for miners to set hashrate at the optimal level. Targeted incentivizes miners to home in on a target hashrate interval. This is achieved by imposing a floor to the miner’s block reward when hashrate is below target and imposing a ceiling on the miner’s block reward when hashrate is above target. The ceiling is progressively lowered and the floor is progressively raised for so long as hashrate is outside the target interval. Targeted achieves aggregate and distributional monetary neutrality. Aggregate monetary neutrality means that the UTXO spending potential of the network is the same under Targeted and Nakamoto at each block. Distributional monetary neutrality means that the UTXO spending potential of addresses are adjusted to offset increases and decreases in the block reward in proportion to their UTXO value.
Identifying the optimal hashrate interval entails some difficulty. Three of the key inputs; estimating hashrate, ASIC efficiency and electricity cost, are, in principle, solvable. But quantifying the other two inputs; the costs of $CO^{2}$ emissions and network security, are elusive. To implement Targeted at the present time will require the formation of a consensus among network participants concerning those quantifications.